### v1

# Towards Unbiased and Robust Spatio-Temporal Scene Graph Generation and Anticipation

Abstract
Spatio-Temporal Scene Graphs (STSGs) provide a concise and expressive representation of dynamic scenes by modelling objects and their evolving relationships over time. However, real-world visual relationships often exhibit a long-tailed distribution, causing existing methods for tasks like Video Scene Graph Generation (VidSGG) and Scene Graph Anticipation (SGA) to produce biased scene graphs. To this end, we propose ImparTail, a novel training framework that leverages curriculum learning and loss masking to mitigate bias in the generation and anticipation of spatio-temporal scene graphs. Our approach gradually decreases the dominance of the head relationship classes during training and focuses more on tail classes, leading to more balanced training. Furthermore, we introduce two new tasks—Robust Spatio-Temporal Scene Graph Generation and Robust Scene Graph Anticipation—designed to evaluate the robustness of STSG models against distribution shifts. Extensive experiments on the Action Genome dataset demonstrate that our framework significantly enhances the unbiased performance and robustness of STSG models compared to existing methods.
Figure 1: (a) Long Tailed Distribution. Predicates in Spatio-Temporal Scene Graph (STSG) datasets exhibit a long-tailed distribution; one such example is the Action Genome [17] dataset, whose distribution is described at the top right. (b) Tasks. We focus on two STSG tasks, including Video Scene Graph Generation (VidSGG) on the left and Scene Graph Anticipation (SGA) on the right. VidSGG entails the identification of fine-grained relationships between the objects observed in the video, such as (Person, looking_at, Paper Notebook) and (Person, not_looking_at, Paper Notebook) in respective frames to the left. (c) Conventional Learning. Due to the inherent long-tailed distribution of these datasets, models learnt using the conventional approaches focus more on the head classes and perform poorly on the tail classes as illustrated using the prediction scores of STTran [7] on contacting and attention relationships (refer middle row). (d) Unbiased Learning. To alleviate the dominance of head classes during training, in unbiased learning, we focus more on the tail classes, ensuring that the learnt models exhibit significantly better performance in predicting both head and tail classes (refer bottom row).
1 Introduction
As AI systems evolve to interact with complex, real-world environments, accurate, adaptive modeling becomes essential. In particular, fields such as autonomous driving [23], human-robot interaction [29], and security surveillance [49] require predictive understanding that goes beyond conventional computer vision. Spatio-Temporal Scene Graphs (STSGs) represent a powerful approach that can address this challenge; at a high level, STSGs are structured graphs where nodes correspond to objects and edges capture the evolving relationships between them over time [17, 23]. Unlike static scene graphs, STSGs offer nuanced video understanding, enabling autonomous systems to interpret interactions in real-time, anticipate changes, and make informed decisions [41].
Yet, constructing effective STSGs that mirror the intricacies of the dynamic, ever-changing real world remains an unsolved problem. In particular, real-world relationships follow a highly imbalanced, long-tailed distribution, with a handful of head classes appearing frequently [44, 2]. In contrast, the vast majority of so-called tail classes are rare but essential for a detailed understanding. This imbalance often skews the learned models towards common relationships, yielding an incomplete, incorrect, unsafe, and distorted interpretation of the scene [42]. Addressing this challenge is vital: without a nuanced, detailed, and balanced representation, AI systems risk overlooking the subtle yet critical interpretations that define complex scenes.
To tackle this challenge, we introduce ImparTail, an innovative framework designed to overcome the pervasive long-tail distribution bias in STSG generation. Via a novel curriculum learning strategy [12], ImparTail shifts attention from common head classes to tail classes, promoting a balanced representation that accurately captures all relationship types, not just the dominant ones. Additionally, ImparTail employs loss masking techniques [27, 2] that amplify the influence of rare classes during training, ensuring a more equitable learning process. This controlled shift towards tail classes allows models to address data distribution disparities without compromising accuracy on head classes.
Real-world data poses an additional challenge beyond bias: distributional shifts. Factors such as lighting, occlusions, and background changes often degrade model performance, affecting the reliability of scene graphs in real-world applications [54, 40]. Conventional models struggle to generalize across these shifts, limiting their effectiveness in unpredictable conditions. To address this, we introduce two new evaluation tasks—Robust Spatio-Temporal Scene Graph Generation and Robust Scene Graph Anticipation—that simulate realistic distributional shifts to assess model resilience under deployment-like conditions.
We rigorously assess the performance of ImparTail on the Action Genome dataset [17], a key benchmark for video understanding. Our results show that ImparTail achieves high mean recall scores. Additionally, ImparTail exhibits robust performance amid distributional shifts, demonstrating that it is well-suited for building robust real-world systems. Thus, this paper makes the following contributions:
  • ImparTail. We introduce a novel, effective, unbiased learning framework for generating and anticipating Spatio-Temporal Scene Graphs by leveraging Partial Stochastic Gradient Descent with Curriculum Learning [12, 28]. Our method is especially suited to address the challenges posed by the long-tailed distribution of real-world relationships.
  • Robustness Evaluation. To assess the resilience of our model against real-world challenges, we systematically induce various corruptions in the input data [42, 20]. We thoroughly estimate the robustness and generalization capabilities by evaluating the model’s performance during the inference phase under these distribution shifts.
  • Empirical Validation. We conduct extensive experiments on the Action Genome [17] benchmark dataset and compare it with existing state-of-the-art methods to empirically validate the efficacy of the proposed approach.
2 Related Work
Tasks. Image Scene Graph Generation (ImgSGG) focuses on representing static visual data—such as 2D and 3D images—as spatial graphs where objects are nodes and their relationships are edges. The field gained significant traction with the foundational Visual Genome project [24], which advanced research in 2D ImgSGG. Building upon this work, Kim et al. [23] extended the task to static 3D scene data by incorporating both RGB and depth information. Object interactions over time provide richer contextual information for dynamic visual content like videos. Transforming this content into structured Spatio-Temporal Scene Graphs (STSGs)—with nodes representing objects and edges capturing temporal relationships—is known as Video Scene Graph Generation (VidSGG). Research in VidSGG has concentrated on improving representation learning through advanced object-centric architectures such as STTran [7] and RelFormer [39]. Shifting the focus from the identification and generation of scene graphs, Peddi et al. [35] recently introduced the Scene Graph Anticipation (SGA) task, which aims to predict STSGs for future frames. Alongside these developments, foundation models have propelled advancements in various ImgSGG task variants [6, 22, 55, 53, 26].
Unbiased Learning. TEMPURA [30] and FlCoDe[21] address the challenges posed by long-tailed datasets, such as those found in Action Genome [17] and VidVRD [38] and propose methods for unbiased VidSGG. Specifically, FloCoDe [21] mitigates bias by emphasizing temporal consistency and correcting the imbalanced distribution of visual relationships. Similarly, TEMPURA [30] addresses biases in relationship prediction with memory-guided training to generate balanced relationship representations and applies a Gaussian Mixture Model to reduce predictive uncertainty. Note. To the best of our knowledge, we are the first to systematically investigate model biases in the SGA task and assess the robustness of both VidSGG and SGA models.
Long-Tail Learning. Long-tailed distributions with a few dominant classes (head classes) often overshadow a more significant number of underrepresented ones (tail classes). This class imbalance typically results in models that perform well on head classes but struggle to generalize to tail classes. To mitigate them, the research community has made significant strides in four directions, which include (a) Cost-Sensitive Learning [14, 45, 8, 1, 47, 32, 45, 50, 51], (b) Mixtures-Of-Experts [48] , (c) Resampling Techniques [4, 9, 52], and (d) Specialized Architectures [52, 54].
Curriculum Learning (CL) is a training methodology that structures training by presenting simpler examples first and progressively introducing more complex ones. This approach aims to enhance learning efficiency by aligning the difficulty of training data with the model’s learning capacity at each stage [25, 11, 12, 13]. Despite its potential benefits, implementing CL presents significant challenges. A primary obstacle is the non-trivial task of distinguishing between easy and hard training samples. This differentiation often requires additional mechanisms, such as auxiliary neural networks acting as teachers or specialized algorithms. Difficulty measures can be predefined based on certain heuristics [16] or learned automatically during the training process [25, 18, 46, 19, 36, 12, 28]. Alongside difficulty assessment, a scheduling strategy is essential to determine when and how to introduce more challenging data [11].
Figure 2: Overview of ImparTail (a) Pipeline. The forward pass of ImparTail begins with an ORPU, where initial object proposals are generated for each observed frame. These object representations are then fed to STPUs designed to construct spatio-temporal context-aware relationship representations of interacting objects. ImparTail applied to both tasks VidSGG and SGA remains mostly the same, with an additional LDPU unit added for SGA to anticipate relationship representations for future frames. These observed (for VidSGG)/ anticipated (for SGA) relationship representations are then decoded to construct STSGs. (b) Conventional Loss. Previous approaches estimated loss for all relationship predicates (head and tail classes). (c) Masked Loss. With the inherent long-tailed nature of the STSG datasets, conventional training results in biased VidSGG and SGA models. Thus, to de-bias the training and learn an unbiased model, in ImparTail, instead of estimating loss for all relationship predicates, we estimate a masked loss, where we selectively mask the labels corresponding to dominant head classes and void their contribution in learning. (d) Curriculum Learning. In ImparTail, we introduce a curriculum-based approach for masking relationship predicate labels during training. At each iteration, we adjust the selection of masked predicates to balance the class distribution progressively. As illustrated, initially, the model trains on the original, long-tailed distribution. As training advances, we systematically mask predicate labels from the head classes, gradually shifting the distribution toward uniformity.
3 Notation & Background
We borrow the notation from SceneSayer[35] and briefly present it below. Let $\textstyle\mathcal{D}$ correspond to the ground truth annotations of an STSG dataset. Consider a video segment $\textstyle V_{1}^{T}$ from a video $\textstyle V$, modelled as a sequence of frames at discrete time steps $\textstyle t=1,2,\ldots,\bar{T}$:
$\textstyle V_{1}^{T}=\{I^{1},I^{2},\ldots,I^{T}\}.$
Here, $\textstyle T$ denotes the total number of frames in the segment, which may vary between videos. In each frame $\textstyle I^{t}$, we represent the scene using a scene graph that captures the visual objects and their pairwise relationships. Let $\textstyle O^{t}=\{o_{1}^{t},o_{2}^{t},\ldots,o_{N(t)}^{t}\}$ be the set of observed objects in frame $\textstyle I^{t}$, where $\textstyle N(t)$ is the number of objects in that frame. Each object $\textstyle o_{k}^{t}$ is characterized by: A bounding box $\textstyle\mathbf{b}_{k}^{t}\in[0,1]^{4}$, representing its spatial location and A category label $\textstyle\mathbf{c}_{k}^{t}\in\mathcal{C}$, where $\textstyle\mathcal{C}$ is the set of all object categories. To describe relationships between objects, we define $\textstyle\mathcal{P}$ as the set of all predicate classes representing possible spatio-temporal relationships. For any pair of objects $\textstyle(o_{i}^{t},o_{j}^{t})$, multiple predicates $\textstyle\{p_{ijk}^{t}\}$ may be associated, where each predicate $\textstyle p_{ijk}^{t}\in\mathcal{P}$. A relationship instance is then defined as the triplet:
$\textstyle r_{ijk}^{t}=(o_{i}^{t},p_{ijk}^{t},o_{j}^{t}).$
Thus, the scene graph $\textstyle\mathcal{G}^{t}$ for frame $\textstyle t$ consists of all such relationship triplets:
$\textstyle\mathcal{G}^{t}=\{r_{ijk}^{t}\}_{i,j,k}.$
For each observed object $\textstyle o_{i}^{t}$ and each object pair $\textstyle(o_{i}^{t},o_{j}^{t})$, we define probability distributions over object categories and predicate classes, respectively: $\textstyle\hat{\mathbf{c}}_{i}^{t}\in[0,1]^{|\mathcal{C}|}$ is the probability distribution over object categories for $\textstyle o_{i}^{t}$ and $\textstyle\hat{\mathbf{p}}_{ij}^{t}\in[0,1]^{|\mathcal{P}|}$ is the probability distribution over predicate classes for the pair $\textstyle(o_{i}^{t},o_{j}^{t})$. These distributions are normalized, satisfying:
$\textstyle\sum_{k}\hat{c}_{ik}^{t}=1,\quad\sum_{k}\hat{p}_{ijk}^{t}=1.$
We distinguish and clarify the objectives of Video Scene Graph Generation (VidSGG) and Scene Graph Anticipation (SGA) as follows: (a) VidSGG. The primary goal of VidSGG is to construct a sequence of scene graphs $\textstyle\{\mathcal{G}^{t}\}_{t=1}^{T}$ corresponding to the observed video segment $\textstyle V_{1}^{T}=\{I^{t}\}_{t=1}^{T}$. This involves: Identifying objects $\textstyle\{o_{k}^{t}\}_{k=1}^{N(t)}$ within each frame $\textstyle I^{t}$ and Determining all possible pairwise relationships $\textstyle\{r_{ijk}^{t}\}_{i,j,k}$ among the detected objects. (b) SGA. The main objective of SGA is to generate scene graphs $\textstyle\{\mathcal{G}^{t}\}_{t=T+1}^{T+H}$ for future frames $\textstyle V_{T+1}^{T+H}=\{I^{t}\}_{t=T+1}^{T+H}$ of the video, using information from the observed segment $\textstyle V_{1}^{T}$. Here, $\textstyle H$ denotes the anticipation horizon [^1] Following the definition and assumptions from SceneSayer [35], we assume the continuity of observed objects in future frames—that is, $\textstyle\{o_{i}^{t}\}_{i=1}^{N(t)}=\{o_{i}^{T}\}_{i=1}^{N(T)}$ for all $\textstyle t>T$—and focus on predicting the evolution of their relationships $\textstyle\{\{r_{ijk}^{t}\}_{t=T+1}^{T+H}\}_{i,j,k}$..
Graph Building Strategies
Building upon the literature in Video Scene Graph Generation (VidSGG) and Scene Graph Anticipation (SGA), we implement three distinct strategies for constructing scene graphs:
  1. 1.
    With Constraint Strategy: This strategy imposes a unique interaction constraint between every pair of objects within a scene. Specifically, for any two objects $\textstyle(o_{i}^{t},o_{j}^{t})$, there exists exactly one relationship predicate $\textstyle p_{ij}^{t}$ (for each relationship category) that describes their interaction. We include these relationship triplets $\textstyle\{r_{ij}^{t}\}_{ij}$ in the scene graph $\textstyle\mathcal{G}^{t}$.
  2. 2.
    No Constraint Strategy: This approach allows for a more expressive graph structure by permitting multiple relationship predicates between any pair of interacting objects $\textstyle(o_{i}^{t},o_{j}^{t})$. We incorporate all predicted relationship triplets $\textstyle\{r_{ijk}^{t}\}_{i,j,k}$ into the scene graph $\textstyle\mathcal{G}^{t}$.
  3. 3.
    Semi-Constraint Strategy: This strategy strikes a balance by permitting multiple relationships between object pairs $\textstyle(o_{i}^{t},o_{j}^{t})$ only if their confidence scores exceed a predefined threshold. Specifically, we add relationship triplets $\textstyle\{r_{ijk}^{t}\}_{i,j,k}$ to the scene graph $\textstyle\mathcal{G}^{t}$ when the confidence score satisfies $\textstyle\hat{p}_{ijk}^{t}>\theta_{k}$.
4 Technical Approach
We propose ImparTail, a framework for unbiased Video Scene Graph Generation (VidSGG) and Scene Graph Anticipation (SGA). In ImparTail, we adopt an object-centric relationship representation processing pipeline as illustrated in Fig 2. Thus, it can be easily extended to de-bias any VidSGG or SGA method built on an object-centric relationship representation processing pipeline [^2] Most VidSGG [7, 10, 30, 21] and SGA [35] approaches share a similar object-centric relationship representation processing pipeline. We further note that recent SOTA VidSGG methods still rely on the object-centric pipelines introduced by STTran [7] and DSGDetr [10].. ImparTail comprises four key components. (I) Object Representation Processing Unit (ORPU): This module generates representations for objects detected in video frames. (II) Spatio-Temporal Context Processing Unit (STPU): That constructs object-centric relationship representations tailored for two tasks: observed relationships for VidSGG and anticipated relationships for SGA. (III) Relationship Predicate Decoders: These decoders predict the relationship predicate labels corresponding to the observed or anticipated relationship representations [^3] Note that ORPU, STPU and Predicate Decoders can be adapted from any VidSGG or SGA methods utilizing an object-centric pipeline.. (IV) Curriculum-Guided Masked Loss: We implement a curriculum-based masking approach that selectively omits the contribution of specific relationship predicate labels during each training iteration. Thus, this method allows the model to progressively balance the class distribution by concentrating on underrepresented classes.
The following sections provide a comprehensive overview of the key components and training methodologies of ImparTail. Specifically, In Sec. 4.1, we outline a detailed explanation of ORPU and STPU, including their respective inputs and outputs within a single training iteration. In Sec. 4.2, we provide a detailed description of the proposed Curriculum-Guided Mask Generation Strategy, and in Sec. 4.3, we present the final change imparted in the loss function for both VidSGG and SGA tasks.
Figure 3: (a) Robustness Evaluation Pipeline: We outline a methodology to evaluate the robustness of trained VidSGG and SGA models under input distribution shifts. Specifically, we introduce systematic corruptions into the frames of test videos, which are then fed as input to the trained models. (b) Corrupted Frames: We illustrate the resultant frames obtained by inducing various categories of corruptions.
4.1 ImparTail Pipeline
Both ORPU and STPU are designed to be compatible with any VidSGG or SGA method. As most existing approaches use similar processing pipelines, we outline the essential architectural components required for our framework to effectively de-bias their methods [^4] Current methods in the literature typically incorporate at least one additional component beyond those specified here. Our experiments directly integrated the ORPU and STPU units from the selected methods.
(I) Object Representation Processing Unit (ORPU): We first extract visual features of observed objects using a pre-trained object detector. Specifically, we extract visual features ($\textstyle\{\mathbf{v}^{t}_{i}\}_{i=1}^{N(t)}$), bounding boxes ($\textstyle\{\mathbf{b}^{t}_{i}\}_{i=1}^{N(t)}$), and object category distributions ($\textstyle\{\hat{\mathbf{c}}^{t}_{i}\}_{i=1}^{N(t)}$), for object proposals $\textstyle\{o^{t}_{i}\}_{i=1}^{N(t)}$ in the observed frames. We use a transformer encoder [43] to integrate temporal information and generate consistent object representations. It is given by:
$\mathbf{V}_{i}^{(n)}=\operatorname{ObjectEncoder}\left(\mathbb{Q}=\mathbb{K}=\mathbb{V}=\mathbf{V}_{i}^{(n-1)}\right)$ (1)
(II) Spatio-Temporal Context Processing Unit (STPU): To learn spatio-temporal context-aware relationship representations, we first construct representations for interacting objects, then refine spatial context with an encoder, and finally use a transformer encoder to integrate spatio-temporal context. Specifically, we define the relationship representation for each observed frame $\textstyle I^{t}$ as $\textstyle\mathbf{z}_{ij}^{t}$, where $\textstyle\mathbf{Z}^{t}$ is formed by stacking all relationship features $\textstyle\{\mathbf{z}_{ij}^{t}\}_{ij}$ corresponding to the objects within the frame. This stacked representation $\textstyle\mathbf{Z}^{t}$ is then passed through the transformer encoder to produce spatial-context-aware representations. Next, we construct the matrix $\textstyle\mathbf{Z}_{ij}$ by stacking the relationship representations $\textstyle\{\mathbf{z}_{ij}^{t}\}_{t=1}^{T}$ over all observed frames. This matrix $\textstyle\mathbf{Z}_{ij}$ is then passed through an encoder, which aggregates temporal information across the stacked relationship representations:
${[\mathbf{Z}^{t}}]^{(n)}=\operatorname{SpatialEncoder}\left(\mathbb{Q}=\mathbb{K}=\mathbb{V}=[{\mathbf{Z}^{t}}]^{(n-1)}\right)\\ \mathbf{Z}_{ij}^{(n)}=\operatorname{TemporalEncoder}\left(\mathbf{Q}=\mathbf{K}=\mathbf{V}=\mathbf{Z}_{ij}^{(n)}\right)$ (2)
(III) Relationship Predicate Decoders: We utilize a two-layer multi-layer perceptron (MLP) to decode the output relationship representations from STPU. We note that for SGA task, following SceneSayer [35], we employ two relationship predicate decoders. One for the observed relationship representations and the other for the anticipated representations. They can be described as follows:
$\hat{\mathbf{p}}^{t}_{ij}=\operatorname{PredClassifier}_{\text{observed}}\left(\mathbf{z}^{t}_{ij}\right),\forall t\in[1,T]\\ \hat{\mathbf{p}}^{t}_{ij}=\operatorname{PredClassifier}_{\text{anticipated}}\left(\mathbf{z}^{t}_{ij}\right),\forall t\in[T+1,\bar{T}]$ (3)
4.2 Curriculum-Guided Mask Generation
Devising curriculum learning approaches for any problem necessitates precise, problem-specific definitions of three key components [28]: (a) the ability to rank tasks from easy to hard, (b) the capability to determine mastery over a specific task, and (c) a procedure that periodically integrates easier tasks with more challenging ones. In this vein, ImparTail employs a simple yet significantly effective curriculum learning method to generate masks for relationship predicate labels during training VidSGG and SGA models.
Banking on the transformers’ prowess in sequential processing tasks, we design the curriculum-guided mask generation process for VidSGG and SGA tasks using simple heuristics as follows: (a) Ranking: We assume that more data associated with a particular relationship predicate label indicates a better ability to predict that class. Therefore, we rank the hardness of relationship predicate learning based on the quantity of data available for them. (b) Mastery: We employ recall as the metric to assess mastery over predicting a specific relationship predicate class, while mean recall evaluates overall performance across all classes. (c) Progressive Mixing: We implement a distributional strategy that progressively balances the class distribution, gradually moving toward an equilibrium where all classes are uniformly represented. The above discussion yields Alg 1, where, as described, in each iteration, we progressively increase the number of masks over dominant relationship predicate labels, voiding their contribution.
4.3 Masked Loss a.k.a Partial SGD
During training, at each epoch $\textstyle e$, as described above, curriculum-guided $\textstyle\operatorname{MaskGenerator}$ provides us with a set of masks $\textstyle\mathcal{M}^{(e)}$ for all the video segments in the dataset $\textstyle\mathcal{D}$. Now, let $\textstyle\mathcal{L}_{p^{t}_{ij}}$ represent any loss defined over a single relationship predicate. Upon application of the mask, the resultant masked loss $\textstyle\mathscr{L}_{p^{t}_{ij}}$ is given by:
$\mathscr{L}_{p^{t}_{ij}}=\mathbf{m}^{t}_{ij}*\mathcal{L}_{p^{t}_{ij}}=\begin{cases}\mathcal{L}_{p^{t}_{ij}}&\text{if }\mathbf{m}^{t}_{ij}=1\\ 0&\text{if }\mathbf{m}^{t}_{ij}=0\end{cases}$ (4)
Input: Epoch: $\textstyle e$, Sampling ratio: $\textstyle\mathcal{S}$, Dataset Annotations: $\textstyle\mathcal{D}$
Output: Ground Truth Annotation Masks: $\textstyle\mathcal{M}$
1
2Compute total annotations $\textstyle N=\sum_{v,f}|\mathcal{D}[v][f]|$
3
4*** Determine Target Counts ***
[…]
6
7*** Curriculum-based sampling probabilities $\textstyle Prob[rel]$**
8
9*** For an equally weighted multinomial distribution ***
10
11Set $\textstyle Prob[rel]=\frac{1}{|\mathcal{P}|}$
12 Sample target counts $\textstyle Tar[rel]$ from Multinomial distribution: $\textstyle Tar[rel]\sim\operatorname{Multinomial}(N_{\text{target}},Prob[rel])$
13
14*** Randomly sample instances of relationship predicates in the dataset based on the target counts for each relationship predicate and construct the filtered dataset ***
15
16Construct filtered dataset $\textstyle\mathcal{F}$
17
18*** Reconstruct Filtered Ground Truth Annotations ***
19 Initialize empty mask list $\textstyle\mathcal{M}$
20 foreach Video $\textstyle v$ in $\textstyle\mathcal{G}$ do
21       Initialize video mask $\textstyle\mathcal{M}_{v}$
22       foreach Frame $\textstyle f$ in $\textstyle v$ do
23             Initialize frame mask $\textstyle\mathcal{M}_{f}$
24             foreach Object $\textstyle o$ in $\textstyle f$ do
25                   Initialize object mask $\textstyle\mathcal{M}_{o}$
26                   Original relations $\textstyle\mathcal{R}_{o}$ from $\textstyle\mathcal{D}$
27                   Filtered relations $\textstyle\mathcal{F}_{o}$ from $\textstyle\mathcal{F}_{r}$
28                   foreach Relation $\textstyle rel$ in $\textstyle\mathcal{R}_{o}$ do
29                         if $\textstyle rel\in\mathcal{F}_{o}[v][f]$ then
30                               Set mask value $\textstyle\mathcal{M}_{o}[rel]=1$
31
32                        else
33                               Set mask value $\textstyle\mathcal{M}_{o}[rel]=0$
34
35
36                  Add $\textstyle\mathcal{M}_{o}$ to $\textstyle\mathcal{M}_{f}$
37
38            Add $\textstyle\mathcal{M}_{f}$ to $\textstyle\mathcal{M}_{v}$
39
40      Add $\textstyle\mathcal{M}_{v}$ to $\textstyle\mathcal{M}$
41
42
Algorithm 1 Mask Generation
VidSGG.
models trained using ImparTail, comprise the combination of (1) Object Classification Loss on the object representations and (2) Masked Predicate Classification Loss on the observed relationship representations as their training objective [^5] Object and predicate losses above can be adapted from any method..Thus, the resultant objective is:
$\underbrace{\mathcal{L}_{i}=\sum_{t=1}^{\bar{T}}\mathcal{L}_{i}^{t}}_{\text{(1)}};\underbrace{\mathscr{L}_{\text{gen}}=\sum_{t=1}^{\bar{T}}\mathscr{L}_{\text{gen}}^{t},\mathscr{L}_{\text{gen}}^{t}=\sum_{ij}\mathbf{m}^{t}_{ij}*\mathcal{L}_{p^{t}_{ij}}}_{\text{{Masked} Predicate Classification Loss (2)}}\quad\\ \mathcal{L}=\sum_{t=1}^{\bar{T}}\left(\lambda_{1}\mathscr{L}_{\text{gen}}^{t}+\lambda_{2}\sum_{i}\mathcal{L}_{i}^{t}\right)$
$\underbrace{\mathcal{L}_{i}=\sum_{t=1}^{\bar{T}}\mathcal{L}_{i}^{t},}_{\text{(1)}};\underbrace{\mathscr{L}_{\text{gen}}=\sum_{t=1}^{\bar{T}}\mathscr{L}_{\text{gen}}^{t},\mathscr{L}_{\text{gen}}^{t}=\sum_{ij}\mathbf{m}^{t}_{ij}*\mathcal{L}_{p^{t}_{ij}}}_{\text{{Masked} Observed Predicate Classification Loss (2)}}\\ \underbrace{\mathscr{L}_{\text{ant}}^{(1:T)}=\sum_{t=T+1}^{\min(T+H,\bar{T})}\mathscr{L}_{\text{ant}}^{t},\quad\mathscr{L}_{\text{ant}}^{t}=\sum_{ij}\mathbf{m}^{t}_{ij}*\mathcal{L}_{p^{t}_{ij}}}_{\text{{Masked} Anticipated Predicate Classification Loss (3)}}\\ \mathcal{L}=\sum_{t=1}^{\bar{T}}\left(\lambda_{1}\mathscr{L}_{\text{gen}}^{t}+\lambda_{2}\sum_{i}\mathcal{L}_{i}^{t}\right)+\\ \sum_{T=3}^{\bar{T}-1}\left(\lambda_{3}\mathscr{L}_{\text{ant}}^{(1:T)}+\lambda_{4}\mathcal{L}_{\text{boxes}}^{(1:T)}+\lambda_{5}\mathcal{L}_{\text{recon}}^{(1:T)}\right)$ (5)
SGA.
models trained using ImparTail, comprise a combination of masked losses over two types of relationship representations: (a) Observed Relationship Representations. and (b) Anticipated Relationship Representations. and other method specific losses. Thus, the objective when ImparTail is employed to SceneSayer [35] includes (1) Object Classification Loss, (2) Masked Predicate Classification Loss over observed relationship representations, (3) Masked Predicate Classification Loss over anticipated relationship representations, (4) Bounding Box Regression Loss and (5) Reconstruction Loss and is given in Eq 5.
4.4 Robustness Evaluation
To evaluate the robustness of VidSGG and SGA models against distribution shifts caused by input corruptions, we introduce two tasks: Robust Video Scene Graph Generation and Robust Scene Graph Anticipation. Fig 3 illustrates the evaluation pipeline established to assess the trained models under these distribution shifts induced by input corruptions. As depicted in the figure, we employ a standard set of corruptions commonly used in adversarial robustness research.
Table 1: Mean Recall Results for VidSGG.
Mode Method With Constraint No Constraint Semi Constraint
mR@10 mR@20 mR@50 mR@10 mR@20 mR@50 mR@10 mR@20 mR@50
SGDET STTran [7] 8.0 16.6 19.3 19.3 26.9 35.6 7.7 18.2 30.4
   +ImparTail (Ours) 9.4 (+17.5%) 21.5 (+29.5%) 25.9 (+34.2%) 23.5 (+21.8%) 33.6 (+24.9%) 43.8 (+23.0%) 8.6 (+11.7%) 21.8 (+19.8%) 38.3 (+26.0%)
DSGDetr [10] 6.7 14.7 19.1 23.3 29.8 36.0 6.5 16.0 30.4
   +ImparTail (Ours) 7.5 (+11.9%) 17.8 (+21.1%) 23.7 (+24.1%) 27.5 (+18.0%) 35.2 (+18.1%) 43.3 (+20.3%) 7.3 (+12.3%) 18.4 (+15.0%) 36.6 (+20.4%)
SGCLS STTran [7] 25.0 27.5 27.6 38.8 47.1 59.9 29.5 39.9 40.9
   +ImparTail (Ours) 32.3 (+29.2%) 36.2 (+31.5%) 36.2 (+31.2%) 47.4 (+22.2%) 57.5 (+22.1%) 66.6 (+11.2%) 36.2 (+22.7%) 50.5 (+26.6%) 52.2 (+27.6%)
DSGDetr [10] 25.6 28.1 28.1 39.9 49.4 64.6 30.1 40.6 41.6
   +ImparTail (Ours) 32.2 (+25.8%) 36.0 (+28.1%) 36.0 (+28.1%) 48.8 (+22.3%) 59.6 (+20.6%) 70.1 (+8.5%) 36.8 (+22.3%) 52.4 (+29.1%) 54.9 (+32.0%)
PREDCLS STTran [7] 30.5 34.7 34.8 45.7 63.4 80.5 36.6 51.8 53.8
   +ImparTail (Ours) 44.0 (+44.3%) 52.7 (+51.9%) 52.9 (+52.0%) 65.5 (+43.3%) 82.0 (+29.3%) 93.0 (+15.5%) 47.7 (+30.3%) 69.7 (+34.6%) 73.4 (+36.4%)
DSGDetr [10] 31.5 36.1 36.2 45.6 64.4 80.5 36.5 52.5 55.2
   +ImparTail (Ours) 41.0 (+30.2%) 48.1 (+33.2%) 48.2 (+33.1%) 59.4 (+30.3%) 76.2 (+18.3%) 89.8 (+11.6%) 43.9 (+20.3%) 65.4 (+24.6%) 69.8 (+26.4%)
Table 2: Mean Recall Results for SGA.
$\textstyle\mathcal{F}$ Method AGS PGAGS GAGS
With Constraint No Constraint With Constraint No Constraint With Constraint No Constraint
@10 @20 @50 @10 @20 @50 @10 @20 @50 @10 @20 @50 @10 @20 @50 @10 @20 @50
0.5 STTran++ [35] 7.9 16.4 18.4 13.9 21.3 38.5 14.3 15.8 15.8 20.9 32.5 50.1 17.8 20.9 21.0 25.2 39.4 63.5
   +ImparTail (Ours) 9.3 18.7 20.9 12.1 20.0 39.6 20.2 22.0 22.0 24.0 34.9 50.5 19.9 22.7 22.8 23.7 39.2 64.1
DSGDetr++ [35] 7.4 13.4 14.6 11.8 18.2 36.1 15.0 16.3 16.3 19.9 32.3 50.6 17.1 20.0 20.0 23.2 37.3 62.9
   +ImparTail (Ours) 8.9 17.0 18.6 13.1 21.6 39.6 18.6 20.1 20.1 22.6 35.2 52.5 21.2 24.5 24.6 28.2 42.2 64.9
SceneSayerODE [35] 5.8 12.6 16.9 14.0 22.3 36.5 11.2 12.8 12.8 16.9 26.3 45.7 17.5 20.7 20.9 24.9 38.0 61.8
   +ImparTail (Ours) 6.8 16.1 22.0 15.6 24.8 39.7 14.5 16.4 16.4 22.7 33.6 49.7 19.3 23.2 23.5 26.5 40.9 63.2
SceneSayerSDE [35] 6.4 13.7 18.3 15.4 23.7 38.7 15.2 17.5 17.5 22.9 34.3 51.0 18.2 21.7 21.8 25.0 39.0 62.7
   +ImparTail (Ours) 7.4 19.1 27.7 21.8 31.4 45.4 15.7 17.9 17.9 23.6 34.3 50.6 17.8 21.2 21.4 27.0 40.7 63.6
0.7 STTran++ [35] 9.1 18.2 20.2 15.7 23.7 41.9 17.2 18.6 18.6 25.3 38.3 56.1 21.9 25.0 25.0 31.2 47.0 75.4
   +ImparTail (Ours) 10.9 21.9 24.1 14.0 23.2 43.7 21.0 22.7 22.7 28.0 41.7 57.1 25.8 29.1 29.1 31.1 49.2 76.5
DSGDetr++ [35] 8.4 14.8 16.0 13.2 20.0 38.8 18.1 19.4 19.4 24.8 39.5 57.3 20.8 23.8 23.8 28.6 46.1 73.8
   +ImparTail (Ours) 10.5 19.5 21.2 14.9 24.8 43.9 20.6 21.8 21.8 26.3 41.0 58.1 28.3 32.5 32.5 31.4 49.7 75.7
SceneSayerODE [35] 6.7 14.0 18.5 16.4 24.9 40.5 13.6 15.1 15.1 20.5 32.4 52.8 20.7 24.0 24.0 29.8 45.2 72.0
   +ImparTail (Ours) 6.8 13.9 18.2 17.5 25.8 41.1 22.2 25.6 25.7 30.7 43.9 55.9 23.2 27.5 27.5 31.7 49.9 73.8
SceneSayerSDE [35] 7.1 14.6 19.3 17.3 26.1 42.5 17.9 19.9 19.9 27.0 40.2 57.2 21.0 24.6 24.6 30.2 45.4 72.8
   +ImparTail (Ours) 8.6 21.3 29.3 25.6 35.1 50.0 25.9 30.0 30.1 35.5 48.2 58.5 20.9 24.4 24.4 31.6 47.9 73.4
Table 3: Robustness Evaluation Results for VidSGG.
$\textstyle\mathcal{S}$ Mode Corruption Method With Constraint No Constraint Semi Constraint
mR@10 mR@20 mR@50 R@10 R@20 R@50 mR@10 mR@20 mR@50 mR@10 mR@20 mR@50
3 SGCLS Gaussian Noise DSGDetr [10] 9.6 10.3 10.3 20.9 25.4 26.8 15.7 19.4 23.4 11.4 15.3 15.7
   +ImparTail (Ours) 13.7 (+42.7%) 14.9 (+44.7%) 15.0 (+45.6%) 21.0 (+0.5%) 27.3 (+7.5%) 30.1 (+12.3%) 20.8 (+32.5%) 25.4 (+30.9%) 29.1 (+24.4%) 15.6 (+36.8%) 21.5 (+40.5%) 22.2 (+41.4%)
Shot Noise DSGDetr [10] 10.0 10.8 10.8 21.9 26.6 28.1 16.6 20.4 26.5 12.1 16.3 16.6
   +ImparTail (Ours) 15.0 (+50.0%) 16.5 (+52.8%) 16.5 (+52.8%) 22.6 (+3.2%) 29.4 (+10.5%) 32.6 (+16.0%) 22.4 (+34.9%) 27.3 (+33.8%) 31.0 (+17.0%) 17.2 (+42.1%) 23.3 (+42.9%) 24.2 (+45.8%)
Impulse Noise DSGDetr [10] 8.7 9.4 9.5 19.2 23.4 24.7 14.4 17.7 23.4 10.4 14.2 14.6
   +ImparTail (Ours) 12.4 (+42.5%) 13.7 (+45.7%) 13.7 (+44.2%) 19.4 (+1.0%) 25.3 (+8.1%) 27.9 (+13.0%) 18.5 (+28.5%) 22.8 (+28.8%) 26.2 (+12.0%) 13.9 (+33.7%) 19.2 (+35.2%) 20.1 (+37.7%)
Speckle Noise DSGDetr [10] 12.6 13.7 13.7 26.2 32.2 34.2 20.3 25.0 32.2 15.0 20.7 21.2
   +ImparTail (Ours) 17.6 (+39.7%) 19.3 (+40.9%) 19.3 (+40.9%) 26.4 (+0.8%) 34.5 (+7.1%) 38.4 (+12.3%) 27.1 (+33.5%) 32.6 (+30.4%) 37.3 (+15.8%) 20.4 (+36.0%) 27.8 (+34.3%) 28.9 (+36.3%)
Fog DSGDetr [10] 22.6 24.9 24.9 47.8 58.4 62.0 35.5 43.4 54.6 26.6 36.1 37.2
   +ImparTail (Ours) 28.3 (+25.2%) 31.8 (+27.7%) 31.9 (+28.1%) 42.6 (-10.9%) 56.0 (-4.1%) 62.1 (+0.2%) 43.8 (+23.4%) 53.0 (+22.1%) 61.7 (+13.0%) 31.8 (+19.5%) 45.7 (+26.6%) 48.2 (+29.6%)
Frost DSGDetr [10] 16.7 18.5 18.5 34.5 42.3 45.1 26.8 33.0 40.1 19.6 26.8 27.7
   +ImparTail (Ours) 22.4 (+34.1%) 25.0 (+35.1%) 25.1 (+35.7%) 31.9 (-7.5%) 42.0 (-0.7%) 47.1 (+4.4%) 34.9 (+30.2%) 42.3 (+28.2%) 48.2 (+20.2%) 25.9 (+32.1%) 36.5 (+36.2%) 38.4 (+38.6%)
Brightness DSGDetr [10] 23.6 25.7 25.7 50.8 61.9 65.5 36.8 45.4 57.5 27.6 37.5 38.6
   +ImparTail (Ours) 29.8 (+26.3%) 33.2 (+29.2%) 33.2 (+29.2%) 45.5 (-10.4%) 59.9 (-3.2%) 66.1 (+0.9%) 45.0 (+22.3%) 55.4 (+22.0%) 65.3 (+13.6%) 33.7 (+22.1%) 48.1 (+28.3%) 50.6 (+31.1%)
Sun Glare DSGDetr [10] 12.1 13.2 13.2 26.3 32.5 34.7 19.3 24.4 30.2 14.2 19.2 19.6
   +ImparTail (Ours) 17.3 (+43.0%) 19.4 (+47.0%) 19.4 (+47.0%) 25.8 (-1.9%) 34.3 (+5.5%) 38.5 (+11.0%) 26.6 (+37.8%) 32.2 (+32.0%) 37.3 (+23.5%) 19.4 (+36.6%) 27.6 (+43.7%) 29.0 (+48.0%)
5 PREDCLS Gaussian Noise STTran [7] 20.0 22.3 22.4 64.2 87.6 99.0 31.4 52.5 79.7 26.0 36.6 38.5
   +ImparTail (Ours) 37.6 (+88.0%) 43.8 (+96.4%) 43.9 (+96.0%) 62.5 (-2.6%) 84.6 (-3.4%) 99.0 (0.0%) 57.5 (+83.1%) 77.7 (+48.0%) 92.7 (+16.3%) 42.2 (+62.3%) 60.0 (+63.9%) 62.9 (+63.4%)
Shot Noise STTran [7] 20.3 22.8 22.9 64.6 88.0 99.0 32.1 53.4 79.9 26.4 37.1 39.0
   +ImparTail (Ours) 37.5 (+84.7%) 43.7 (+91.7%) 43.8 (+91.3%) 62.4 (-3.4%) 84.6 (-3.9%) 99.0 (0.0%) 57.1 (+77.9%) 77.9 (+45.9%) 92.8 (+16.1%) 42.2 (+59.8%) 60.5 (+63.1%) 63.6 (+63.1%)
Impulse Noise STTran [7] 20.3 22.8 22.9 64.6 88.0 99.0 31.9 52.9 79.8 26.3 37.2 39.1
   +ImparTail (Ours) 37.7 (+85.7%) 43.8 (+92.1%) 43.9 (+91.7%) 62.4 (-3.4%) 84.7 (-3.8%) 99.0 (0.0%) 57.4 (+79.9%) 77.7 (+46.9%) 92.6 (+16.0%) 42.0 (+59.7%) 60.1 (+61.6%) 63.1 (+61.4%)
Speckle Noise STTran [7] 23.1 26.1 26.2 66.8 89.6 99.1 37.1 57.8 80.3 30.2 43.1 45.3
   +ImparTail (Ours) 38.8 (+68.0%) 45.5 (+74.3%) 45.5 (+73.7%) 60.2 (-9.9%) 82.9 (-7.5%) 98.4 (-0.7%) 59.4 (+60.1%) 77.8 (+34.6%) 92.5 (+15.2%) 43.6 (+44.4%) 60.7 (+40.8%) 63.1 (+39.3%)
Fog STTran [7] 26.5 30.2 30.3 70.2 91.1 99.1 41.6 61.0 80.5 33.2 46.8 48.7
   +ImparTail (Ours) 42.6 (+60.8%) 50.9 (+68.5%) 51.1 (+68.6%) 64.8 (-7.7%) 86.3 (-5.3%) 98.8 (-0.3%) 63.8 (+53.4%) 80.2 (+31.5%) 92.7 (+15.2%) 46.2 (+39.2%) 65.5 (+40.0%) 68.2 (+40.0%)
Frost STTran [7] 25.6 29.2 29.2 69.4 90.7 99.1 41.0 60.9 80.5 32.7 46.1 48.0
   +ImparTail (Ours) 41.0 (+60.2%) 49.0 (+67.8%) 49.2 (+68.5%) 62.2 (-10.4%) 84.3 (-7.1%) 98.5 (-0.6%) 62.5 (+52.4%) 78.6 (+29.1%) 92.7 (+15.2%) 45.1 (+37.9%) 62.9 (+36.4%) 65.1 (+35.6%)
Brightness STTran [7] 28.2 32.0 32.1 71.3 91.6 99.2 42.8 62.0 80.4 34.5 49.0 51.2
   +ImparTail (Ours) 42.3 (+50.0%) 50.4 (+57.5%) 50.5 (+57.3%) 65.9 (-7.6%) 87.2 (-4.8%) 98.9 (-0.3%) 64.0 (+49.5%) 80.8 (+30.3%) 92.8 (+15.4%) 46.9 (+35.9%) 67.8 (+38.4%) 71.0 (+38.7%)
Sun Glare STTran [7] 22.5 25.1 25.2 66.7 89.9 99.1 36.7 56.7 80.0 28.9 40.5 42.3
   +ImparTail (Ours) 40.2 (+78.7%) 47.5 (+89.2%) 47.7 (+89.3%) 57.9 (-13.2%) 81.7 (-9.1%) 98.0 (-1.1%) 60.3 (+64.3%) 77.5 (+36.7%) 92.7 (+15.9%) 43.3 (+49.8%) 59.3 (+46.4%) 61.0 (+44.2%)
Table 4: Robustness Evaluation Results for SGA.
$\textstyle\mathcal{F}$ Mode Corruption Method With Constraint
mR@10 mR@20 mR@50
0.5 PGAGS Gaussian Noise STTran+ [35] 7.9 8.4 8.4
   +ImparTail (Ours) 5.1 (-35.4%) 5.4 (-35.7%) 5.4 (-35.7%)
DSGDetr+ [35] 5.5 5.8 5.8
   +ImparTail (Ours) 7.5 (+36.4%) 8.0 (+37.9%) 8.0 (+37.9%)
STTran++ [35] 5.9 6.4 6.4
   +ImparTail (Ours) 9.4 (+59.3%) 10.2 (+59.4%) 10.2 (+59.4%)
DSGDetr++ [35] 5.7 6.1 6.1
   +ImparTail (Ours) 8.3 (+45.6%) 8.8 (+44.3%) 8.8 (+44.3%)
Fog STTran+ [35] 9.8 10.6 10.6
   +ImparTail (Ours) 11.4 (+16.3%) 12.2 (+15.1%) 12.2 (+15.1%)
DSGDetr+ [35] 10.8 12.0 12.0
   +ImparTail (Ours) 15.2 (+40.7%) 16.6 (+38.3%) 16.6 (+38.3%)
STTran++ [35] 11.7 12.9 12.9
   +ImparTail (Ours) 15.8 (+35.0%) 17.4 (+34.9%) 17.4 (+34.9%)
DSGDetr++ [35] 12.7 13.8 13.8
   +ImparTail (Ours) 15.8 (+24.4%) 17.4 (+26.1%) 17.4 (+26.1%)
Frost STTran+ [35] 8.2 9.0 9.0
   +ImparTail (Ours) 8.3 (+1.2%) 8.7 (-3.3%) 8.7 (-3.3%)
DSGDetr+ [35] 9.0 9.9 9.9
   +ImparTail (Ours) 12.4 (+37.8%) 13.3 (+34.3%) 13.3 (+34.3%)
STTran++ [35] 9.6 10.5 10.5
   +ImparTail (Ours) 13.8 (+43.7%) 15.3 (+45.7%) 15.3 (+45.7%)
DSGDetr++ [35] 9.9 10.8 10.8
   +ImparTail (Ours) 13.2 (+33.3%) 14.4 (+33.3%) 14.4 (+33.3%)
Brightness STTran+ [35] 11.0 12.1 12.1
   +ImparTail (Ours) 11.7 (+6.4%) 12.5 (+3.3%) 12.5 (+3.3%)
DSGDetr+ [35] 12.2 13.5 13.5
   +ImparTail (Ours) 15.9 (+30.3%) 17.2 (+27.4%) 17.2 (+27.4%)
STTran++ [35] 12.8 14.1 14.1
   +ImparTail (Ours) 17.7 (+38.3%) 19.3 (+36.9%) 19.3 (+36.9%)
DSGDetr++ [35] 13.6 14.7 14.7
   +ImparTail (Ours) 16.6 (+22.1%) 18.1 (+23.1%) 18.1 (+23.1%)
5 Experiments
We apply ImparTail to generate and anticipate STSGs on the Action Genome dataset [17]. We prepare the dataset by following the pre-processing step described in [7, 35]. The dataset includes 35 object classes and 25 relationship classes. These 25 relationship classes are grouped into three categories, namely: (1) Attention Relations comprise relationship classes which primarily describe the attention of the subject towards the object, (2) Spatial Relations comprise relationship classes that describe the spatial relationship between the subject and objects, and (3) Contacting Relations comprises relationship classes that indicate different ways the subject contacts the object  [^6] Remark. The dataset predominantly features videos with a single individual interacting with objects across diverse real-world settings. Within the Action Genome framework, a subject-object pair can exhibit multiple types of spatial and contact relationships..
Evaluation Metric. We evaluate our models using the standard Recall@K and meanRecall@K metrics, where $\textstyle K$ takes values within the set {10, 20, 50}. The Recall@K metric measures the model’s capability to predict the relationships between observed objects in future frames. The long-tailed distribution of relationships in the training set [30] can generate biased scene graphs, causing frequently occurring relationships to dominate the Recall@K metrics. To address this imbalance, we utilize the mean recall metric introduced in [5], a more balanced metric that scores the model’s generalisation to all predictive classes.
In the following sections, we will assess the performance of ImparTail on four tasks, namely: (a) VidSGG, (b) SGA, (c) RobustVidSGG and (d) RobustSGA. For each task, we first select a few methods, train and evaluate them as per the instructions, then adapt them to ImparTail framework, re-train and evaluate them. For VidSGG and Robust VidSGG tasks, we selected STTran [7] and DSGDetr [10] as our baselines and for the SGA task, we selected STTran+ [35], STTran++ [35], DSGDetr+ [35], DSGDetr++ [35], SceneSayerODE [35], SceneSayerSDE [35] as our baselines.
Settings.
We follow [7] and use three different settings to evaluate models for the VidSGG task. Each setting has a different level of information provided as input to the models. (a) Scene Graph Detection (SGDET): We only provide frames of the video as input, (b) Scene Graph Classification (SGCLS): Along with frames, bounding box information is provided as input to the model and (c) Predicate Classification (PREDCLS): We provide frames, bounding boxes and object labels as input to the model. Similarly, for the SGA task, we follow SceneSayer [35] and use three different settings to evaluate models. (a) Action Genome Scenes (AGS): In AGS, the model’s input is limited to raw frames of the video. (b) Partially Grounded Action Genome Scenes (PGAGS): In this intermediate setting, along with raw video frames, we additionally input the model with precise bounding box information of active interacting objects observed in the scene. (c) Grounded Action Genome Scenes (GAGS): In this model takes precise bounding box information and the categories of the observed interacting objects as input.
5.1 Video Scene Graph Generation.
We began by training and evaluating the selected baseline models. Subsequently, we adapted these models to incorporate ImparTail, followed by retraining and evaluation. Our experiments encompassed three VidSGG modes SGDET, SGCLS, and PREDCLS and were reported under three distinct settings: (1) With Constraint, (2) No Constraint, and (3) Semi Constraint. As demonstrated in Tab 1, models trained using ImparTail consistently outperformed baselines.
[…]
5.2 Scene Graph Anticipation.
To train and evaluate the SGA models, we adopted a methodology similar to that of VidSGG. Initially, we trained and assessed the selected baseline models. Subsequently, we integrated these models into the ImparTail framework, followed by retraining and re-evaluating their performance. As demonstrated in Table 2, Models trained using ImparTail consistently outperformed their baseline variants.
[…]
5.3 Robustness Evaluation.
We evaluated VidSGG and SGA models trained with ImparTail and those trained using traditional approaches across 16 input corruption scenarios. Our experiments revealed that introducing noise to the inputs generally led to a decline in performance. However, models trained with ImparTail exhibited a less severe performance drop than baseline models. For example, in the PREDCLS mode under the No Constraint setting with Gaussian Noise for STTran, the baseline model’s mR@10 decreased from 45.7 to 31.4. In contrast, the ImparTail-trained model only dropped from 65.5 to 57.5. To maintain brevity, only a subset of the results are reported in the main paper, and we defer additional results to the supplementary text.
Insights: (1) We noticed that the drop in performance of both VidSGG and SGA models trained using ImparTail is lower than the models trained with traditional approaches. (2) as illustrated in Tables 3, 4, our models performed significantly better than the baseline variants, (3) We also observed that models trained using ImparTail not only exhibited consistent strong mean recall scores but also outperformed others in recall metrics sometimes (refer Tab 3).
6 Conclusion
In this paper, we tackled the pervasive challenge of long-tailed distribution bias in Spatio-Temporal Scene Graph (STSG) generation, often leading to models that overlook rare but critical relationships in dynamic scenes. We introduced ImparTail, a framework that employs curriculum learning and loss masking techniques to shift the learning focus toward tail classes without sacrificing performance on head classes. To further evaluate the practical applicability of our model, we proposed two new tasks: Robust Spatio-Temporal Scene Graph Generation and Robust Scene Graph Anticipation. These tasks simulate real-world distributional shifts, allowing us to assess the robustness of STSG. Through extensive experiments on the Action Genome, ImparTail demonstrated superior performance in mean recall scores and maintained robustness against various input corruptions. Future work includes developing robust learning techniques [15, 33, 3] for both VidSGG and SGA tasks and exploring the application of unbiased VidSGG and SGA in long-tailed regimes such as error recognition [34, 37], action anticipation [31] etc.